# TrimCaching: Parameter-sharing Edge Caching for AI Model Downloading

### v1

@@ Abstract
Edge AI model caching, edge computing, edge intelligence, 6G.
@@ I Introduction
In the era of Artificial Intelligence of Things (AIoI), there is an ongoing trend for pushing Artificial Intelligence (AI) services from the cloud to local devices due to the stringent latency and privacy requirements of diverse AI-empowered applications [2, 3, 4, 5]. For instance, autonomous driving must perceive the surrounding environments in a timely fashion [6, 7, 8]; the emerging large language models (LLMs) often involve privacy-sensitive personal data for human-machine interactions [9, 10]. In line with this evolution, next-generation mobile networks aim to support rapid model downloading for mobile users to realize “on-device AI” with better latency and privacy guarantees. Specifically, due to the sheer number of AI applications and the ever-growing size of AI models (e.g., Google’s on-device LLM Gemini Nano-2 with 3.25 billion parameters), it is impractical for users to store every AI model locally. By ensuring efficient model downloading, 6th generation (6G) mobile networks enable mobile users to enjoy versatile on-device AI services with low latency without exhausting their data storage capacities.
@@
Conventional model downloading features the delivery of AI models from the remote cloud to local devices, which is infeasible for latency-sensitive services. For example, as specified by 3GPP 5G technical specifications, both autonomous vehicles and robots may need to accomplish AI model downloading within 1 second [11]. Considering the large-sized models and the limited bandwidth of cloud centers, the required fast model downloading may only be achieved by placing AI models in edge networks, leading to a paradigm called “edge model caching”. However, unlike cloud centers, an edge server can only store a limited number of popular models under the storage capacities of edge servers. To enhance model downloading performance, one fundamental research problem, therefore, is model placement, e.g., how can we optimally place AI models on edge servers to improve the cache hit ratio under the latency requirements of various AI services?
@@
In this paper, we identify an AI model placement problem by observing parameter sharing among AI models. Shared parameters pervade AI models nowadays, which can be exploited to store them more efficiently at the network edge. A broad range of AI models, such as convolutional neural networks (CNNs) or large language models (LLMs), can share a considerable percentage of parameters since they can be created from the same pre-trained models [12]. For instance, layer freezing is a very classic transfer learning or multi-task learning method because bottom layers in deep neural networks, such as convolution layers, usually share common knowledge reusable for different downstream tasks[13, 14, 15]. In the era of LLMs, parameter-efficient fine-tuning (PEFT) (e.g., LoRA[16]) is highly effective in fine-tuning foundation models to downstream or customized models, where the core idea is to freeze the significant proportion of parameters (e.g., more than 99% in LoRA for LLMs) to save computing and memory resources. Today’s engineers often fine-tune pre-trained models (e.g., pre-trained CNNs on the Pytorch platform or open-source foundation models such as GPT-3) on task-specific datasets to create new services. To demonstrate this phenomenon, Fig. 1 shows that the inference accuracy in ResNet50 only degrades slightly as the number of shared layers grows. Even if the first 90% of trainable layers, up to layer 97, are frozen, the average accuracy degradation is only about 4.7% compared with the full-layer fine-tuning. Considering LLMs, Fig. 2 illustrates that the BLEU scores for the various fine-tuned GPT-2 medium models exhibit slight variation across different frozen parameter ratios. The fine-tuned model retains satisfactory performance, even when up to 99.97% of the parameters are frozen from the pre-trained model. In a nutshell, downstream models can share a significant proportion of parameters from pre-trained models, making storage more efficient.
@@
    We show that the resultant problem is a submodular maximization problem with submodular constraints. After problem mapping, we conclude that there is no polynomial-time algorithm that can solve it with a constant approximation ratio due to the submodular constraints resulting from shared parameter blocks.
@@
    We design a greedy algorithm for the original problem for the general case. Although a constant approximation guarantee cannot be achieved as shown before, the algorithm is still efficient and effective in performance evaluation in the general case.
@@ II Related Work
Content caching in wireless edge networks has attracted significant attention since it brings contents closer to end users, leading to a field called edge caching [26, 27, 28, 29]. Specifically, popular files can be pre-cached in wireless edge servers, such as small base stations [30], thereby enabling content downloading to users with low latency. One fundamental research problem in edge caching is how to place contents on edge servers for enhancing the quality of experience (QoE) of end users, which is called the content placement problem [31, 32]. Due to the overlapping coverage of wireless edge servers (e.g., base stations), users can download files from any of the nearby edge servers with the desired contents. This class of QoE maximization problems has been identified as submodular maximization problems with knapsack constraints [22, 23, 24, 25]. In these schemes, since contents are independent of each other, the storage constraints are naturally knapsack constraints. Simple greedy methods are usually effective in solving such constrained submodular maximization problems with good approximation ratios [22]. In our work, due to the parameter sharing across AI models, the shared parameters only need to be cached once in a server, resulting in the submodularity of storage constraints. For this reason, when adapting the aforementioned content placement strategies, e.g., greedy algorithms [22], to our case, no constant approximation guarantee can be achieved.
In the field of edge intelligence, caching AI models in distributed wireless edge networks can facilitate model delivery for both inference and training [33, 34, 35, 36, 37]. End users can download the required AI models from edge servers, thereby significantly reducing service latency [38, 39]. Analogous to traditional edge caching, these studies focus on optimizing model placement decisions to enhance the QoE of users subject to the storage limitations of edge servers. However, these papers have not considered parameter sharing either. In [40], Wu et al propose a multi-user model downloading method by exploiting the shared parameters in AI models, but their scheme focuses on designing a multicasting scheme rather than model placement. To our best knowledge, our work is the first to define the parameter-sharing model placement problem and develop approximate algorithms for it.
@@ III-A Network Scenario
We consider a typical scenario in wireless edge networks, as shown in Fig. 3. Let the set =$\mathcal{M}=\left\{1,2,\dots,m\dots,M\right\}$ denote the $M$ wireless edge servers (e.g., base stations). All edge servers are interconnected. There are a set ${\mathcal{K}}=\left\{1,2,\dots,k,\dots,K\right\}$ of users covered by these edge servers. The network operator provides a library of AI models =${\mathcal{I}}=\left\{1,2,\dots,i,\dots,I\right\}$, which can be downloaded by users for inference services. The request probability and the QoS constraint on end-to-end (E2E) latency of user $k$ for model $i$ are ${p}_{k,i}$ and $\bar{T}_{k,i}$, respectively. The frequently used notations are summarized in Table I.
@@
  • Downloading from associated edge servers: User $k$ first sends the model request to its associated edge servers. If user $k$ can directly download model $i$ from any edge server $m$ in $\mathcal{M}_{k}$ storing this model and finish the on-device inference under the latency requirement $\bar{T}_{k,i}$, we call it cache hit.
  • Downloading from non-associated edge servers: If no edge server in $\mathcal{M}_{k}$ caches model $i$, the remaining edge servers will provide the model to the user. Suppose that model $i$ is cached on any edge server in $\mathcal{M}\setminus\mathcal{M}_{k}$. Model $i$ will be first transferred to an edge server in $\mathcal{M}_{k}$ for downloading to the user with the lowest latency. We call it cache hit if the E2E latency (including edge-to-edge, edge-to-user and on-device inference latency) meets the latency requirement $\bar{T}_{k,i}$.
@@ III-C Design Objective
The TrimCaching framework aims to place AI models appropriately to serve as many user requests as possible by exploiting shared parameter blocks among models. Since parameter blocks are shared, the storage capacity of edge servers can be leveraged more effectively, which can help edge networks serve more user requests. For cache misses on edge servers, the user can request the models from the cloud center. However, since fetching models from the edge will be much faster than the cloud, we will develop model placement policies to maximize the cache hit ratio on edge servers under latency constraints, which is also a common objective in edge caching [41].
@@ IV-A Problem Formulation
The TrimCaching framework targets to maximize the expected cache hit ratio for users’ model requests by solving the model placement problem under the storage capacity of edge servers and service latency requirements. After normalization, the expected cache hit ratio is
@@
where ${p}_{k,i}$ is the request probability of user $k$ for model $i$. $x_{m,i}$ is the placement of model $i$ on edge server $m$, where $x_{m,i}=1$ means model $i$ is cached on edge server $m$. Besides,
@@
is an indicator function, where ${\mathbb{I}}_{1}\left(m,k,i\right)=1$ if and only if $T_{m,k,i}\leq{\bar{T}}_{k,i}$. $T_{m,k,i}$ is the total latency when edge server $m$ delivers model $i$ to user $k$, including model downloading latency and on-device inference latency. If user $k$ can download model $i$ from its associated edge server, we have
@@
In $U\left({\bf{X}}\right)$, $\prod\limits_{m\in\mathcal{M}}\left(1-x_{m,i}\mathbb{I}_{1}\left(m,k,i\right)\right)=0$ represents user $k$ can obtain the requested model $i$ from any edge server within latency requirement, implying that the request for model $i$ from user $k$ can be served by edge networks. Moreover, we use $Q_{m}$ to denote the storage capacity of edge server $m$. The problem formulation is given as follows.
@@
where $D^{\prime}_{j}$ is the size of parameter block $j$. The data size of parameter $j$ is calculated once per edge server using the product term when it is shared by the corresponding models in $\mathcal{I}$. $11-\prod\limits_{i\in\mathcal{I}_{j}}\left(1-x_{m,i}\right)=$ 1 means that parameter block $j$ is stored exactly once on edge server $m$, because there is at least a model containing block $j$ placed on the server.
Note that our formulation ignores user mobility because the problem is solved based on a “snapshot” of user locations. However, this is commonly adopted in model placement schemes [22]. In practice, our algorithm can conduct model placement decisions by solving the above problem, then re-initiate model placement only when the performance degrades to a certain threshold. Our simulation results will show that our algorithm is resilient to user mobility over time, thus eliminating the need for frequent model replacement that would consume backbone bandwidth.
@@ IV-B Problem Mapping
where ${\bf{V}}_{m}$ is the universal set for edge server $m$. For any element $x_{m,i}\in{\bf{V}}_{m}\setminus{\bf{X}}^{2}_{m}$, the increased occupied storage is $g_{m}\left({\bf{X}}^{1}_{m}\cup\left\{x_{m,i}\right\}\right)-g_{m}\left({\bf{X}}^{1}_{m}\right)$ when we add $x_{m,i}$ into ${\bf{X}}^{1}_{m}$. Since ${\bf{X}}^{1}_{m}\subseteq{\bf{X}}^{2}_{m}$, it is more likely that the parameter blocks contained by model $i$ have already been included in models in ${\bf{X}}^{2}_{m}$. Therefore, $g_{m}\left({\bf{X}}^{1}_{m}\cup\left\{x_{m,i}\right\}\right)-g_{m}\left({\bf{X}}^{1}_{m}\right)\geq g_{m}\left({\bf{X}}^{2}_{m}\cup\left\{x_{m,i}\right\}\right)-g_{m}\left({\bf{X}}^{2}_{m}\right)$, which means that the constraint function of $\mathcal{P}1.1$ is a set of submodular functions. ∎
@@
We first introduce two mathematical concepts. First, $\sum\limits_{x\in X}g\left(x\right)\leq b$ is called a knapsack constraint, where $g\left(x\right)$ is the cost function, and $X$ is a finite set. Second, if the output of an algorithm for a maximization problem is at least $\alpha$-times of the optimal result, we can say that this algorithm ensures a constant approximation guarantee. Here $\alpha\in\left[0,1\right]$ and $\alpha$ is a constant.
$\mathcal{P}1.1$ is equal to the following problem $\mathcal{P}1.2$, where $y_{m,j}$ represents the caching status of parameter block $j$ on edge server $m$. $y_{m,j}=1$ means that edge server $m$ stores parameter block $j$. The relationship between $x_{m,i}$ and $y_{m,j}$ are: $x_{m,i}=\prod\limits_{j\in\mathcal{J}_{i}}y_{m,j}$ and $y_{m,j}=1-\prod\limits_{i\in\mathcal{I}_{j}}\left(1-x_{m,i}\right)$. With this relationship, $U\left({\bf{X}}\right)$ is equal to $U\left(\bf{Y}\right)=\frac{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}\left[1-\prod\limits_{m\in\mathcal{M}}\left(1-\prod\limits_{j\in\mathcal{J}_{i}}y_{m,j}{\mathbb{I}}_{1}\left(m,k,i\right)\right)\right]}{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}{p}_{k,i}}$.
@@
$\mathcal{P}1.2$ is a supermodular maximization problem with $M$ knapsack constraints. The proof is omitted here since it is similar to the proof of Proposition 1. It has been proved that the supermodular maximization problem with $M$ knapsack constraints is an NP-hard problem, and there is no approximation algorithm with a constant-approximation guarantee to solve this problem when $M=1$ [44]. Therefore, there is no polynomial-time algorithm with a constant-approximation guarantee for $\mathcal{P}1.2$ and further for $\mathcal{P}1.1$. ∎

### v2

@@ Abstract
Edge AI model caching, edge computing, edge intelligence, 6G, model downloading.
@@ I Introduction
In the era of Artificial Intelligence of Things (AIoI), there is an ongoing trend for pushing Artificial Intelligence (AI) services from the cloud to local devices due to the stringent latency and privacy requirements of diverse AI-empowered applications [2, 3, 4, 5]. For instance, autonomous driving must perceive the surrounding environments in a timely fashion [6, 7, 8]; the emerging large language models (LLMs) often involve privacy-sensitive personal data for human-machine interactions [9, 10]. On the other hand, due to the sheer number of AI applications and the ever-growing size of AI models (e.g., Google’s on-device LLM Gemini Nano-2 with 3.25 billion parameters), it is impractical for users to store every AI model locally. For this reason, next-generation mobile networks aim to support rapid model downloading for mobile users to realize “on-device AI” with better latency and privacy guarantees. By ensuring efficient model downloading, the 6th generation (6G) mobile networks enable mobile users to enjoy versatile on-device AI services with low latency without exhausting their data storage capacities.
@@
Conventional model downloading features the delivery of AI models from the remote cloud to local devices, which is infeasible for latency-sensitive services. For example, as specified by 3GPP 5G technical specifications, both autonomous vehicles and robots may need to accomplish AI model downloading within 1 second [11]. Observing the large-sized models and the limited bandwidth of remote cloud centers, this kind of fast model downloading may only be achieved by placing AI models in edge networks close to users, leading to a paradigm called “edge model caching”. However, unlike cloud centers, an edge server can only store a limited number of popular models under the storage capacities of edge servers. To enhance model downloading performance, one fundamental research problem, therefore, is model placement, e.g., how can we optimally place AI models on edge servers to maximize the cache hit ratio for model downloading under the latency requirements of various AI services?
@@
In this paper, we identify an AI model placement problem by observing parameter sharing among AI models. Shared parameters pervade AI models nowadays, which can be exploited to store them more efficiently at the network edge. A broad range of AI models, such as convolutional neural networks (CNNs) or large LLMs, can share a considerable percentage of parameters since they can be created from the same pre-trained models [12]. For instance, layer freezing is a very classic transfer learning or multi-task learning method because bottom layers in deep neural networks, such as convolution layers, usually share common knowledge reusable for different downstream tasks[13, 14, 15]. In the era of LLMs, parameter-efficient fine-tuning (PEFT) (e.g., LoRA[16]) is highly effective in fine-tuning foundation models to downstream or customized models, where the core idea is to freeze the significant proportion of parameters (e.g., more than 99% in LoRA for LLMs) to save computing and memory resources. Today’s engineers often fine-tune pre-trained models (e.g., pre-trained CNNs on the Pytorch platform or open-source foundation models such as GPT-3) on task-specific datasets to create new services. To demonstrate this phenomenon, Fig. 1 shows that the inference accuracy in ResNet50 only degrades slightly as the number of shared layers grows. Even if the first 90% of trainable layers, up to layer 97, are frozen, the average accuracy degradation is only about 4.7% compared with the full-layer fine-tuning. Considering LLMs, Fig. 2 illustrates that the BLEU scores for the various fine-tuned GPT-2 medium models exhibit slight variation across different frozen parameter ratios. The fine-tuned model retains satisfactory performance, even when up to 99.97% of the parameters are frozen from the pre-trained model. In a nutshell, downstream models can share a significant proportion of parameters from pre-trained models, making storage more efficient.
@@
    We show that the resultant problem is a submodular maximization problem with submodular constraints. After problem mapping, we conclude that there is no polynomial-time algorithm to solve it with a constant approximation ratio due to the submodular constraints resulting from shared parameter blocks.
@@
    We design a greedy algorithm for the original problem for the general case. Although a constant approximation guarantee cannot be achieved as alluded to earlier, we show that the algorithm is still efficient and effective through performance evaluation.
@@ II Related Work
Content caching in wireless edge networks has attracted significant attention since it brings contents closer to end users, leading to a field called edge caching [26, 27, 28, 29]. Specifically, popular files can be pre-cached in wireless edge servers, such as small base stations [30], thereby enabling content downloading to users with low latency. One fundamental research problem in edge caching is how to place contents on edge servers for enhancing the quality of experience (QoE) of end users, which is called the content placement problem [31, 32]. Due to the overlapping coverage of wireless edge servers (e.g., base stations), users can download files from any of the nearby edge servers with the desired contents. This class of QoE maximization problems has been identified as submodular maximization problems with knapsack constraints [22, 23, 24, 25]. In these schemes, since contents are independent of each other, the storage constraints are naturally knapsack constraints. Simple greedy methods are usually effective in solving such constrained submodular maximization problems with good approximation ratios [22]. In our work, due to the parameter sharing across AI models, the shared parameters only need to be cached once in a server, resulting in the submodularity of storage constraints. For this reason, when adapting the aforementioned content placement strategies, e.g., greedy algorithms [22], to our case, no theoretical constant approximation guarantee can be achieved.
In the field of edge intelligence, caching AI models in distributed wireless edge networks can facilitate model delivery for both inference and training [33, 34, 35, 36, 37]. End users can download the required AI models from edge servers, thereby significantly reducing service latency [38, 39]. Analogous to traditional edge caching, these studies focus on optimizing model placement decisions to enhance the QoE of users subject to the storage limitations of edge servers. However, these papers have not considered parameter sharing either. In [40], Wu et al. propose a multi-user model downloading method by exploiting the shared parameters in AI models, but their scheme focuses on designing a multicasting scheme rather than model placement. To our best knowledge, our work is the first to define the parameter-sharing model placement problem, identify its mathematical properties, and develop the corresponding solution approaches.
@@ III-A Network Scenario
We consider a typical scenario in wireless edge networks, as shown in Fig. 3. Let the set =$\mathcal{M}=\left\{1,2,\dots,m\dots,M\right\}$ denote the $M$ wireless edge servers (e.g., base stations). All edge servers are interconnected. There are a set ${\mathcal{K}}=\left\{1,2,\dots,k,\dots,K\right\}$ of users covered by these edge servers. The network operator provides a library of AI models =${\mathcal{I}}=\left\{1,2,\dots,i,\dots,I\right\}$, which can be downloaded by users for inference services. The request probability and the quality of service (QoS) constraint on end-to-end (E2E) latency of user $k$ for model $i$ are ${p}_{k,i}$ and $\bar{T}_{k,i}$, respectively. The frequently used notations are summarized in Table I.
@@
  • Downloading from associated edge servers: User $k$ first sends the model request to its associated edge servers. If user $k$ can directly download model $i$ from any edge server $m$ in $\mathcal{M}_{k}$ storing this model and finish the on-device inference under the latency requirement $\bar{T}_{k,i}$, a cache hit occurs
  • Downloading from non-associated edge servers: If no edge server in $\mathcal{M}_{k}$ caches model $i$, the remaining edge servers will provide the model to the user. Specifically, model $i$ will be transferred from the edge server caching the model to an edge server in $\mathcal{M}_{k}$ for downloading to the user with the lowest latency. A cache hit occurs if the E2E latency (including edge-to-edge, edge-to-user, and on-device inference latency) meets the latency requirement $\bar{T}_{k,i}$.
@@ III-C Design Objective
The TrimCaching framework aims to judiciously place AI models to serve as many user requests as possible by exploiting shared parameter blocks among models. For cache misses on edge servers, the downloading request can be forwarded to the cloud center for fetching the model. However, since downloading models from the cloud can be much slower, our policies aim to maximize the cache hit ratio, defined as the probability of successfully downloading the AI models from edge servers within latency constraints, which is also a common design objective in edge caching [41].
@@ IV-A Problem Formulation
The TrimCaching framework aims to maximize the expected cache hit ratio for users’ model requests by addressing the model placement problem under the storage capacity of edge servers and service latency requirements. The expected cache hit ratio is
@@
where ${p}_{k,i}$ is the request probability of user $k$ for model $i$. $x_{m,i}$ is the placement of model $i$ on edge server $m$, where $x_{m,i}=1$ indicates that model $i$ is cached on edge server $m$. $\prod\limits_{m\in\mathcal{M}}\left(1-x_{m,i}\mathbb{I}_{1}\left(m,k,i\right)\right)=0$ implies user $k$ can obtain the requested model $i$ from any edge server within latency requirement, where
@@
is an indicator function, and ${\mathbb{I}}_{1}\left(m,k,i\right)=1$ if and only if $T_{m,k,i}\leq{\bar{T}}_{k,i}$. $T_{m,k,i}$ is the E2E latency when edge server $m$ delivers model $i$ to user $k$, including model downloading latency and on-device inference latency. If user $k$ can download model $i$ from its associated edge server, we have
@@
The problem formulation is given as follows.
@@
where $D^{\prime}_{j}$ is the size of parameter block $j$, and the storage capacity of edge server $m$ is denoted by $Q_{m}$. $11-\prod\limits_{i\in\mathcal{I}_{j}}\left(1-x_{m,i}\right)=$ 1 means that parameter block $j$, if shared by multiple models on edge server $m$, is stored only once, which leads to improved storage efficiency.
Note that our formulation ignores user mobility because the problem is solved based on a “snapshot” of user locations. This is commonly adopted in model placement schemes [22]. In practice, our algorithm can conduct model placement decisions by solving the above problem, then re-initiate model placement when the performance degrades to a certain threshold. Our simulation results will show that our algorithm is resilient to user mobility over time, thus eliminating the need for frequent model replacement that would consume backbone bandwidth.
@@ IV-B Problem Mapping
where ${\bf{V}}_{m}$ is the universal set for edge server $m$. For any element $x_{m,i}\in{\bf{V}}_{m}\setminus{\bf{X}}^{2}_{m}$, the incremental occupied storage is $g_{m}\left({\bf{X}}^{1}_{m}\cup\left\{x_{m,i}\right\}\right)-g_{m}\left({\bf{X}}^{1}_{m}\right)$ when we add $x_{m,i}$ into ${\bf{X}}^{1}_{m}$. Since ${\bf{X}}^{1}_{m}\subseteq{\bf{X}}^{2}_{m}$, it is likely that the parameter blocks contained by model $i$ have already been included in models corresponding to ${\bf{X}}^{2}_{m}$. Therefore, $g_{m}\left({\bf{X}}^{1}_{m}\cup\left\{x_{m,i}\right\}\right)-g_{m}\left({\bf{X}}^{1}_{m}\right)\geq g_{m}\left({\bf{X}}^{2}_{m}\cup\left\{x_{m,i}\right\}\right)-g_{m}\left({\bf{X}}^{2}_{m}\right)$, implying that the constraint function of $\mathcal{P}1.1$ is a set of submodular functions. ∎
@@
We first introduce two mathematical concepts. First, $\sum\limits_{x\in X}g\left(x\right)\leq b$ is called a knapsack constraint, where $g\left(x\right)$ is a cost function and $X$ is a finite set. Second, if the output of an algorithm for a maximization problem is guaranteed to be at least $\alpha$-times of the optimal solution, we can say that this algorithm ensures a constant approximation guarantee. Here $\alpha\in\left[0,1\right]$ and $\alpha$ is a constant.
$\mathcal{P}1.1$ is equal to the following problem $\mathcal{P}1.2$, where $y_{m,j}$ represents the caching status of parameter block $j$ on edge server $m$. $y_{m,j}=1$ indicates that edge server $m$ stores parameter block $j$. The relationship between $x_{m,i}$ and $y_{m,j}$ are: $x_{m,i}=\prod\limits_{j\in\mathcal{J}_{i}}y_{m,j}$ and $y_{m,j}=1-\prod\limits_{i\in\mathcal{I}_{j}}\left(1-x_{m,i}\right)$. With this relationship, $U\left({\bf{X}}\right)$ is transformed into $U\left(\bf{Y}\right)=\frac{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}\left[1-\prod\limits_{m\in\mathcal{M}}\left(1-\prod\limits_{j\in\mathcal{J}_{i}}y_{m,j}{\mathbb{I}}_{1}\left(m,k,i\right)\right)\right]}{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}{p}_{k,i}}$.
@@
$\mathcal{P}1.2$ is a supermodular maximization problem with $M$ knapsack constraints. The proof is omitted here since it is similar to the proof of Proposition 1. It has been shown that the supermodular maximization problem with $M$ knapsack constraints is an NP-hard problem, and there is no approximation algorithm with a constant-approximation guarantee to solve this problem when $M=1$ [44]. Therefore, there is no polynomial-time algorithm with a constant-approximation guarantee for $\mathcal{P}1.2$, which also holds for $\mathcal{P}1.1$. ∎